OpenStudy (anonymous):

Anyone understand what I'm being asked here? Find a function f such that f'(x) = x^3 and the line x+y=0 is tangent to the graph of f.

OpenStudy (anonymous):

so far, as I understand it, they're asking me to find the antiderivative of x^3 whereby f'(x) is x+y=0???

OpenStudy (anonymous):

I mean, this could be anything, right? f(x) = (x^4)/4+2 would work...

OpenStudy (anonymous):

x could then equal -1.

OpenStudy (anonymous):

then x+y=0!

OpenStudy (anonymous):

Do I understand this correctly?